OpenStudy (sloppycanada):

Check my work - How to find the eccentricity of the following eclipse? (x+2)^2/12 + (y-4)^2/9 = 1

OpenStudy (sloppycanada):

The equation eccentricity is sqrt(1-b^2/a^2 Correct? So, this stands to reason that it's - sqrt(1-9^2/12^2)

OpenStudy (sshayer):

\[b^2=a^2\left( 1-e^2 \right),1-e^2=\frac{ b^2 }{ a^2 },e^2=1-\frac{ b^2 }{ a^2 },e=\sqrt{1-\frac{ b^2 }{ a^2 }}\] so you are absolutely correct.

OpenStudy (sshayer):

correction \[a^2=12\] b^2=9 so correct accordingly and show again.